# Prediction of the Most Fire-Sensitive Point in Building Structures with Differentiable Agents for Thermal Simulators

### v1

@@ Acknowledgements
The authors extend their heartfelt gratitude to Dr. M. Eslami, Mr. G. Zhou, Mr. C.M. Perez, Mr. W. Ruan, Mr. K.T. Pang, and Ms. L. Yang, members of STAIRlab, for their valuable discussions and contributions to this research. Special thanks are due to Prof. S.-L. Huang of Tsinghua University and Prof. L. Zheng of MIT for insightful discussions related to this paper. We are deeply appreciative of the Berkeley Research IT team for providing access to the Savio computing cluster, a vital resource for the “STAIRlab” project. This advanced high-performance computing infrastructure enabled data generation and network training, which were critical to the success of this research. Finally, we gratefully acknowledge the financial support provided to STAIRlab through the Taisei Chair in Civil Engineering at UC Berkeley. Both the computational and financial resources were indispensable completing this work.

### v5

@@ Acknowledgements
The authors extend their heartfelt gratitude to Dr. M. Eslami, Mr. G. Zhou, Mr. C.M. Perez, Mr. W. Ruan, Mr. K.T. Pang, and Ms. L. Yang, members of STAIRlab, for valuable discussions related to this research. Special thanks go to Prof. S.-L. Huang (Tsinghua University) and Prof. L. Zheng (MIT) for their insightful discussions that greatly enriched this study. We also deeply appreciate the Berkeley Research IT team for providing access to the Savio computing cluster, a critical resource for the “STAIRlab” project. This high-performance computing infrastructure played a key role in data generation and network training, both of which were essential to the success of this research. Furthermore, we gratefully acknowledge the financial support provided to STAIRlab through the Taisei Chair in Civil Engineering at UC Berkeley. Both the computational and financial resources were indispensable in completing this work. Finally, we sincerely thank the editorial team for their efficient handling of the manuscript. We are particularly indebted to the nine anonymous peer reviewers, whose meticulous evaluations and thoughtful suggestions significantly improved the technical accuracy and overall clarity of this paper.